The mean weight of statistics books in the U.S. is 140 pounds with a standard deviation of 20 lbs.

We sure pack a lot of stuff in these books!!! You select 400 books at random and measured their

weight and found the mean for this sample was 137 lbs. Using a level of significance of 0.05, test the

hypothesis that the true weight is 140 lb.
Answer #1

solution:- Ho: M=140 Hail & 140 L = X-M g e - = -3 137-140 - 20 v 400 m P-value = 0.0027 Since 2=0.05 & P-value=0.0027 P-valu